# Efficient Agent Training for Computer Use

### v1

@@ 4 WindowsAgentArena-V2
Current computer use benchmarks such as WindowsAgentArena and OSWorld often include infeasible tasks, which are inherently impossible to complete due to issues such as deprecated system features or user-generated hallucinated commands Xie et al. (2024). The evaluation metric for these tasks is simply considering a task successful if the action FAIL is output at any point during execution. However, we found such evaluation methods particularly easy to hack, as agents can achieve high scores on infeasible tasks without demonstrating meaningful computer-use capabilities. In contrast, completing a feasible task typically requires the agent to execute actions step-by-step to actually fulfill the task objective, posing a significantly different level of difficulty.
We refer to this phenomenon as infeasible hacking, a vulnerability confirmed by our subsequent experiments (§5.5), in which a weaker model achieved markedly higher scores on infeasible tasks. Since agents receive identical scores for feasible and infeasible tasks, their coexistence undermines benchmark fairness. Additionally, given that current computer use agents’ abilities are far from optimal, we argue it is presently more valuable to focus on enhancing agents’ performance on feasible tasks. Therefore, as a temporary solution in WindowsAgentArena-V2, we removed all infeasible tasks to prevent infeasible hacking.
@@ 5 Experiment
In this section, we evaluate PC Agent-E through four experiments: main performance assessment on our improved benchmark WindowsAgentArena-V2 (§5.2), train time action scaling to validate the effect of our Trajectory Boost method (§5.3), test time action scaling (§5.4), and cross-platform generalization to Linux environments using OSWorld (§5.5).
@@ 5.1 Setup
We use WindowsAgentArena-V2 (§4) for the main evaluation, as our training data were collected on the Windows system. To test generalization across operating systems, we also report results on OSWorld Xie et al. (2024), a popular benchmark designed on Linux systems.
Baseline
We compare PC Agent-E with various strong baselines. These include the proprietary frontier models Claude 3.7 Sonnet Anthropic (2025a) and Claude 3.7 Sonnet with extended thinking Anthropic (2025b), as well as leading open-source models UI-TARS Qin et al. (2025), UI-TARS-1.5 Team (2025), and Qwen2.5-VL-72B Bai et al. (2025). For UI-TARS-1.5, we evaluated its open-source UI-TARS-1.5-7B version, as its larger counterpart was not accessible.
@@
We train our PC Agent-E model based on the Qwen2.5-VL-72B backbone with 27k data mentioned in §3.5. We set the image resolution to $1280\times 720$ and context length to 8,192 tokens. Further training details can be found in Appendix A.
@@ 5.2 Main Results
As shown in Table 2, our PC Agent-E achieves a remarkable 141% relative improvement over the Qwen2.5-VL-72B baseline on WindowsAgentArena-V2. More significantly, PC Agent-E establishes itself as the SOTA open-source computer use agent on Windows, achieving the best performance on most applications. It is worth highlighting that the Claude 3.7 Sonnet model used to synthesize our training data did not have thinking mode enabled, making our achievement even more substantial as we surpassed the stronger Claude 3.7 Sonnet with extended thinking.
Table 2: Results of successful rate (%) for different models on WindowsAgentArena-V2.
@@
Our result reveals that complex agent capabilities for computer use can be elicited by a remarkably small set of high-quality trajectory data. What lies behind our extreme data efficiency is our high data quality. We attribute this high quality to two critical factors: (1) Real-world task completion, guaranteed through human-annotated task trajectories that reliably record authentic interaction processes and ensure successful task execution; and (2) Diverse action decisions, synthesized by frontier agents model that can generate multiple reasonable actions with rational thought at each step beyond the single solution chosen by human annotators.
@@
To gain deeper insight into the specific capabilities enhanced through our training, we conducted a qualitative analysis by examining 50 task trajectories that Qwen2.5-VL-72B failed to complete but PC Agent-E successfully accomplished, as well as trajectories where both models failed. We categorized the failure patterns into three types, which may occur simultaneously within a single trajectory or even a single step: (1) Knowledge: models may lack specific computer use knowledge. For instance, a model might be unaware of how to enable a particular feature in VLC (a popular media player software), thus failing to complete the task. (2) Planning: models may make incorrect action decisions, such as failing to recognize and recover from previous erroneous actions. (3) Grounding: even when the model’s planning is correct, the executed actions might be inconsistent with its planning process, primarily manifested as mouse-clicking errors.
Analysis indicates that our improvements primarily stem from enhanced long-horizon planning capabilities. After training, PC Agent-E produces noticeably longer thought processes and demonstrates improved reasoning abilities in verification, reflection, and self-correction. Besides, both Qwen2.5-VL-72B and PC Agent-E occasionally fail due to grounding and knowledge issues.
5.3 Train Time Action Scaling
In §3.4, we synthesize 9 alternative action decisions at each step of a human trajectory using the Trajectory Boost method. To validate the effect of this data augmentation method, we vary the number of synthesized actions used during training and evaluate model performance. Let $s$ be the number of synthesized actions used per step, and $s^{\prime}=s+1$ denote the scaling factor of the training data after augmentation. When $s<9$, the $s$ action decisions were randomly selected from the 9 synthesized options. For our final PC Agent-E model, all 9 synthesized actions were included, yielding $s^{\prime}=10$. In contrast, for the model trained solely on human actions, $s^{\prime}=1$. All training data followed the structure detailed in §3.5. As shown in Figure 8(a), model performance significantly scales up with increased action decisions. Compared to training on human trajectories alone (only a 15% relative improvement over the base model), PC Agent-E achieved significantly greater gains, demonstrating the effectiveness of the Trajectory Boost method in enhancing data quality and enabling efficient agent training.
@@ 5.3 Train Time Action Scaling
Figure 8: (a) Performance on WindowsAgentArena-V2 with different training data scaling factor $s^{\prime}$. (b) Performance on WindowsAgentArena-V2 with different max steps before and after training.
5.4 Test Time Action Scaling
We also investigated how the performance of PC Agent-E varies with test time scaling, a topic that has received increasing attention in the research community Wu et al. (2025); Brown et al. (2024); Snell et al. (2024). We evaluated the model’s performance with different numbers of max steps allowed during task completion. As shown in Figure 8(b), the difference in performance between two models increases over time as the agent continues to interact with the computer. With improved long-horizon planning capabilities, PC Agent-E benefits from “action scaling”—the ability to iteratively take more actions to respond to environmental changes, explore for solutions, correct errors, and ultimately solve the task. After training, PC Agent-E can allocate more time and computational resources to computer use tasks to achieve superior results.
@@ 5.5 Cross-Platform Evaluation
We further evaluated our model on the popular OSWorld benchmark to assess cross-platform generalization capabilities. As shown in Table 3, our PC Agent-E model, despite being trained exclusively on Windows data, demonstrates a 34% relative improvement in computer use capabilities on Linux systems as well. These results validate the strong cross-platform generalizability of our efficient agent training methodology.
Table 3: Successful Rate (%) on OSWorld for 30-step evaluation.
@@
We also identified an interesting phenomenon in the experiments, which we designate as infeasible hacking in §4: the weaker Qwen2.5-VL-72B model paradoxically achieved markedly better performance on infeasible tasks. This observation suggests that current infeasible task evaluations do not accurately reflect a computer use agent’s capabilities. Future research may focus on designing stronger evaluation criteria for infeasible tasks, including assessing an agent’s rationale when declaring a task as impossible, to minimize exploitative behaviors.
6 Conclusion and Discussion
In this work, we introduced PC Agent-E, an efficient agent training framework for computer use. With just 312 human-annotated trajectories augmented through our Trajectory Boost method, PC Agent-E achieved a 141% improvement over the base model and outperformed Claude 3.7 Sonnet with extended thinking on WindowsAgentArena-V2, while demonstrating strong generalizability on OSWorld. Our findings demonstrate that complex agent capabilities for computer use can be elicited by a remarkably small set of high-quality trajectory data.
Looking forward, with the significant breakthroughs of RL in reasoning models, researchers have recently begun applying RL to long-horizon tasks OpenAI (2025a, b, c) and achieving encouraging progress. However, several studies Guo et al. (2025); Yue et al. (2025) indicate that RL’s effectiveness is closely tied to the capabilities of the foundation models. Currently, the computer use capabilities of foundation models remain far from sufficient, even in state-of-the-art proprietary models employing RL OpenAI (2025a). We have reason to believe that collecting computer use data for pre-training and post-training remains crucial for achieving truly intelligent digital world agents. The future direction will likely involve the collaborative development of RL and SFT, complementing and enhancing each other.
7 Limitations
There are multiple potential limitations in this work, primarily due to resource constraints. First, we only used Claude 3.7 Sonnet for data synthesis rather than more advanced models like OpenAI Operator OpenAI (2025a), though our impressive results suggest stronger models would likely yield further improvements. Second, we only fine-tuned one model, Qwen2.5-VL-72B, as our representative backbone. Third, we did not explore how increasing trajectory count beyond our current dataset might affect performance, yet our results with extremely limited trajectories already demonstrate remarkable data efficiency.
@@ Appendix A Training Details
Appendix B Prompts
B.1 Thought Completion
@@
You are a helpful computer use agent designed to complete tasks on a computer. Your goal is to recreate your thought process behind a specific action. You will be provided with: 1. The task you are attempting to complete. 2. A history of the steps you have already performed (up to 50, if any; none if it was the first action). 3. The specific action you chose to take. 4. The name of the element you clicked (if you clicked on an element). It might be too general or vague, you have to decied what to click based on the screenshot. 5. A screenshot of the computer screen at the moment you decided to take the action. 6. The red marks on the screenshot indicate the position of the click or drag action. To formulate your thought process, consider: 1. What do you observe on the screen? Consider your task and previous action when you analyzing current screenshot. 2. Evaluate your previous action (if applicable): - Did it achieve the intended effect? If not, identify possible reasons (e.g., misclick, inactive element). Some typical examples for ineffective action: - misclick in an empty space - ineffective opening some elements without double click - ineffective type text/ press key because of inactivated input box - Did the result align with your previous plan, or did something unexpected happen? Some typical examples for ineffective action: - misclick in a wrong element - forget to clear existing text in input bar 3. Based on your action history, assess your progress toward completing the overall task. 4. Consider if youre exploring how to finish the task because of failed attempts in history steps. Present your thought process as a clear, natural first-person narrative that explains your reasoning at that moment. Important requirements: 1. **DO NOT** mention the red marks in your response. These marks were **added after the fact** to indicate the position of your click or drag actions, and they were not on the screen when you made the decision. **DO NOT** mention "red box", "red square", "red circle", or "red arrow" in your response. 2. Write as if you are thinking in real-time before taking the action. Do not include post-action evaluation or hindsight. The task you are attempting to complete: {task_description} Your performing history: {history_str} The specific action you chose to perform: {action}
B.2 Trajectory Boost
Table 5: Trajectory Boost Prompt
Prompt for Trajectory Boost
You are a helpful assistant who can help users complete computer tasks, with **full permission** to make any operations on the users computer. The operating system is windows. Based on the provided current state, you need to suggest the next action to complete the task. Do not try to complete the entire task in one step. Break it down into smaller steps, and at each step you will get a new state to interact with. IMPORTANT: You must strictly adhere to the following rules: 1. Choose ONLY ONE action from the list below for each response, DO NOT perform more than one action per step. 2. Follow the exact syntax format for the selected action, DO NOT create or use any actions other than those listed. 3. Once the task is completed, output action finish. Valid actions: 1. click (x, y) click the element at the position (x, y) on the screen 2. right click (x, y) right click the element at the position (x, y) on the screen 3. double click (x, y) double click the element at the position (x, y) on the screen 4. drag from (x1, y1) to (x2, y2) drag the element from position (x1, y1) to (x2, y2). 5. scroll (x) scroll the screen vertically with pixel offset x. Positive values of x: scroll up, negative values of x: scroll down. 6. press key: key_content press the key key_content on the keyboard. 7. hotkey (key1, key2) press the hotkey composed of key1 and key2. 8. hotkey (key1, key2, key3) press the hotkey composed of key1, key2, and key3. 9. type text: text_content type content text_content on the keyboard. Note that before typing text, you need to ensure the text box or input field is active/focused first. If the text box is not yet activated, you should first click on it to activate it, and then use type text in a separate step. 10. wait wait for some time, usually for the system to respond, screen to refresh, advertisement to finish.
[…]
B.3 PC Agent-E scaffold
Table 6: PC Agent-E scaffold Prompt
Prompt for PC Agent-E scaffold
You are a helpful assistant who can help users complete computer tasks, with **full permission** to make any operations on the users computer. Based on the provided current state, you need to suggest the next action to complete the task. Do not try to complete the entire task in one step. Break it down into smaller steps, and at each step you will get a new state to interact with. IMPORTANT: You must strictly adhere to the following rules: 1. Choose ONLY ONE action from the list below for each response, DO NOT perform more than one action per step. 2. Follow the exact syntax format for the selected action, DO NOT create or use any actions other than those listed. 3. Once the task is completed, output action finish. Valid actions: 1. click (x, y) click the element at the position (x, y) on the screen 2. right click (x, y) right click the element at the position (x, y) on the screen 3. double click (x, y) double click the element at the position (x, y) on the screen 4. drag from (x1, y1) to (x2, y2) drag the element from position (x1, y1) to (x2, y2). 5. scroll (x) scroll the screen vertically with pixel offset x. Positive values of x: scroll up, negative values of x: scroll down. 6. press key: key_content press the key key_content on the keyboard. 7. hotkey (key1, key2) press the hotkey composed of key1 and key2. 8. hotkey (key1, key2, key3) press the hotkey composed of key1, key2, and key3. 9. type text: text_content type content text_content on the keyboard. 10. wait wait for some time, usually for the system to respond, screen to refresh, advertisement to finish. 11. finish indicating that the task has been completed. 12. fail indicating that the task has failed, of this task is infeasible because not enough information is provided.
Response Format: {Your thought process} Action: {The specific action you choose to take} Your task is: {task_description} History of the previous actions and thoughts you have done to reach the current screen: {history_str} -------------------------------------------- Given the screenshot, whats the next step you will do to help with the task?
Appendix C PC Tracker User Manual
1. Introduction
PC Tracker is a lightweight infrastructure for efficiently collecting large-scale human-computer interaction trajectories. The program runs seamlessly in the background, automatically capturing screenshots and keyboard & mouse activities.
2. Installation
  • Ensure your operating system is Windows.
  • Extract our software package to a location with sufficient disk space (recommended to have more than 3GB of available space for storing recorded data).
3. Quick Start
  • (Optional) Set screen resolution to 16:9.
  • Open the extracted folder and launch main.exe.
4. Instructions
After launching the tracker, you can choose between Task Oriented Mode or Non-Task Oriented Mode for recording.
Task Oriented Mode
This mode is divided into two sub-modes: Given Task and Free Task.
Given Task
In this mode, you will be assigned an uncompleted task each time.
  • Next Task: Click Next Task to get the next task.
  • Previous Task: Click Previous Task to return to the previous task.
  • Bad Task Feedback: If you think the current task is difficult to complete, click Bad Task to discard it permanently. Alternatively, you can start the task and modify its description after completion based on your actual execution.
  • Start Recording: Click Start, and the tracker window will automatically minimize while recording begins.
  • End Task: After completing the task, click Finish to save the record. Or if the task execution fails or you don’t want to record it, click Fail.
  • Modify Task Description: After finishing the task, you can modify the task description based on your actual execution.
Free Task
In this mode, you can freely use the computer and summarize the task description and difficulty yourself.
  • Start Recording: Click Start, and the tracker window will automatically minimize while recording begins.
  • Save and Summarize This Record: Fill in the task description, select difficulty (easy/medium/hard), and click Save to save the record.
  • Discard This Record: Click Discard to discard the record.
Non-Task Oriented Mode
In this mode, you can freely use the computer, with similar methods to start and stop recording as described above.
5. Usage Notes
  • Does not currently support using extended screens.
  • Does not currently support using Chinese input methods.
  • Does not currently support using touchpads.
  • The tracker window is fixed in fullscreen. To support the filtering of tracker-related actions (such as clicking the Start button) in post-processing, the tracker window is fixed in fullscreen. You can reopen the tracker window by clicking to view the task description, then minimize it again, but please do not drag it to display in a non-fullscreen state.
6. Data Privacy
  • After starting recording, your screenshots and keyboard & mouse operations will be automatically recorded. PC Tracker does not record any information from unopened software. If you believe the recording may infringe on your privacy, you can choose to discard the record.
  • Collected data is saved in the ./events folder (hidden by default). Each trajectory includes a Markdown file for easy visualization.
7. FAQ
Does the tracker have networking capabilities?
PC Tracker is completely local, does not support networking, and will not upload your data.
What if my computer screen resolution is not 16:9?
If your screen resolution is not 16:9, it will affect the subsequent unified processing of data. We recommend adjusting your screen resolution to 16:9.
How much space does the collected data occupy?
The specific data size varies. Generally, even with intensive recording operations for 1 hour, it will not generate more than 1GB of data.

### v2

@@ 4 WindowsAgentArena-V2
Current computer use benchmarks such as WindowsAgentArena and OSWorld often include infeasible tasks, which are inherently impossible to complete due to issues such as deprecated system features or user-generated hallucinated commands Xie et al. (2024). The evaluation metric for these tasks is simply considering a task successful if the action FAIL is output at any point during execution. However, we found such evaluation methods particularly easy to hack: an agent can trivially achieve a perfect score on infeasible tasks by always outputting FAIL, without demonstrating any meaningful computer use capabilities. In contrast, completing a feasible task typically requires the agent to execute actions step-by-step to actually fulfill the task objective, posing a significantly different level of difficulty.
We refer to this phenomenon as infeasible hacking, a vulnerability confirmed by our subsequent experiments (§5.6), in which a weaker model achieved markedly higher scores on infeasible tasks. Since agents receive identical scores for feasible and infeasible tasks, their coexistence undermines benchmark fairness. Additionally, given that current computer use agents’ capabilities are far from optimal, we argue it is presently more valuable to focus on enhancing agents’ performance on feasible tasks. Therefore, as a temporary solution in WindowsAgentArena-V2, we removed all infeasible tasks to prevent infeasible hacking.
@@ 5 Experiment
In this section, we conduct extensive experiments to evaluate PC Agent-E and validate our Trajectory Boost method. Our experiments are designed to answer the following key questions:
  1. 1.
    How does PC Agent-E perform against SOTA methods on computer use tasks? (§5.2)
  2. 2.
    How does Trajectory Boost’s data scaling surpass using human demonstrations alone? (§5.3)
  3. 3.
    How does Trajectory Boost differ from and outperform Direct Distillation? (§5.4)
  4. 4.
    How does test-time scaling affect the performance of PC Agent-E? (§5.5)
  5. 5.
    How well does PC Agent-E generalize to unseen environments? (§5.6)
@@ 5.1 Setup
We use WindowsAgentArena-V2 (§4) for the main evaluation, as our training data were collected on the Windows system. We also include results on the original WindowsAgentArena Bonatti et al. (2024) in Appendix B for completeness. To test generalization across operating systems, we report results on OSWorld Xie et al. (2024), another computer use benchmark for Linux systems.
Model Baseline
We compare PC Agent-E with several SOTA models. These include the leading proprietary models Claude 3.7 Sonnet Anthropic (2025a) and Claude 3.7 Sonnet with extended thinking Anthropic (2025b), as well as open-source models UI-TARS Qin et al. (2025), UI-TARS-1.5 Team (2025), and Qwen2.5-VL-72B Bai et al. (2025). We also compared with the popular GPT-4o OpenAI (2024) model.
Method Baseline
We compare our Trajectory Boost method with two alternative training approaches. The first is standard behavior cloning on the 312 human trajectories after thought completion. The second is direct distillation from Claude. We sample 10 end-to-end trajectories from Claude 3.7 Sonnet for each of the 312 tasks. The resulting 3,120 trajectories are then used for training with the identical procedure as PC Agent-E, matching the trajectory number of our method for a fair comparison.
@@
We train our PC Agent-E model based on the Qwen2.5-VL-72B Bai et al. (2025) backbone with 27k data mentioned in §3.5. Experiments on a smaller model Qwen2.5-VL-7B are also included in Appendix C. We set the image resolution to $1280\times 720$ and context length to 8,192 tokens. Further training details can be found in Appendix A.
@@ 5.2 Main Results
As shown in Table 3, PC Agent-E achieves a remarkable 141% relative improvement over the base model Qwen2.5-VL-72B on WindowsAgentArena-V2, even surpassing the strong teacher model Claude 3.7 Sonnet by 10% in relative terms (36.0 vs. 32.6), establishing itself as the SOTA open-source model for Windows computer use. Notably, Claude 3.7 Sonnet used to synthesize our training data did not have the thinking mode enabled, but PC Agent-E achieves performance comparable to the stronger Claude 3.7 Sonnet with extended thinking.
@@
Table 3: Results of success rate (%) for different models on WindowsAgentArena-V2.
@@
To gain deeper insight into the specific capabilities enhanced through our training, we conducted a qualitative analysis by examining 50 trajectories that Qwen2.5-VL-72B failed but PC Agent-E succeeded, as well as trajectories where both models failed. We categorized the failure patterns into three types: (1) Knowledge: models may lack specific computer use knowledge. For instance, a model might not know how to enable a particular feature in VLC (a media player software). (2) Planning: models may make incorrect planning, such as failing to recognize and recover from previous erroneous actions. (3) Grounding: models may execute actions that are inconsistent with their plan, primarily manifested as mouse-clicking errors. We found that our improvements primarily stem from enhanced planning capabilities. After training, PC Agent-E produces noticeably longer thought processes and demonstrates improved reasoning capabilities in verification and self-correction. We did not observe significant improvements in knowledge or grounding capabilities.
5.3 Data Scaling over Human Demonstrations
To validate the effectiveness of our Trajectory Boost method, we investigate the relationship between the scale of synthesized data and model performance. We define data scaling factor, $s$, as the ratio of the total action number used for training to the action number in the original human trajectory. For the model trained exclusively on the human demonstrations, the scaling factor is $s=1$. Our final model, PC Agent-E, was trained using 9 synthesized actions and 1 original human action per step, corresponding to the scaling factor $s=9+1=10$.
As the blue line shown in Figure 7, our results reveal that model performance with the Trajectory Boost method scales significantly with the scaling factor. Compared to training on human trajectories alone, which yields a limited gain (improved from 14.9 to 17.2), PC Agent-E achieves substantially greater performance gains (improved from 14.9 to 36.0). This improvement is primarily driven by the diverse action decisions synthesized from the frontier model with thought processes. This supplements the single human-annotated solution and instills the frontier model’s advanced planning capabilities into our agent, thereby yielding performance that far exceeds training on human trajectories alone.
@@ 5.3 Data Scaling over Human Demonstrations
Figure 7: Performance of Trajectory Boost and Direct Distillation method with different data scaling factor $s$ on WindowsAgentArena-V2.
5.4 Trajectory Boost vs. Direct Distillation
To demonstrate that our method is more than a simple form of distillation, we compare Trajectory Boost against a direct distillation baseline. For this baseline, we directly sample trajectories from our teacher model, Claude 3.7 Sonnet, end-to-end. We also compare with a variant baseline that combines human demonstrations and direct distillation, with details provided in Appendix D.
Superior Performance
As shown in Figure 7, our method significantly outperforms the direct distillation baseline at most of the training data scales (blue line versus orange line). We attribute this to the high quality of our synthesized data. Our method uses human trajectories as a reliable foundation and leverages the frontier model to perform single-step augmentation. This avoids the error accumulation that can occur in end-to-end trajectory distillation.
Exceptional Efficiency
Another significant advantage of our method is time efficiency. The distillation baseline requires deploying Claude in virtual machines and collecting trajectories through online interaction, which makes it resource-intensive and time-consuming. In contrast, our Trajectory Boost method performs offline data synthesis without interacting with the real environment, enabling natural parallelization. Specifically, to collect an equivalent amount (3120 trajectories) of data, the distillation baseline took about 900 hours, while Trajectory Boost required only 3 hours under the same hardware conditions — a drastic 300-fold speedup.
5.5 Test Time Scaling
Figure 8: Test time scaling on WindowsAgentArena-V2.
We also investigate how the performance of PC Agent-E varies with test time scaling, a topic that has received increasing attention in the research community Wu et al. (2025); Snell et al. (2024). We evaluated the model’s performance with different numbers of max steps allowed during task completion. As shown in Figure 8, when the step limit increases from 15 to 30, PC Agent-E effectively leverages the additional compute, with the performance gap over the base model widening as more steps are allowed. This suggests that the improved planning capabilities acquired through training enable PC Agent-E to benefit from longer interaction horizons. We further extend the evaluation to a 50-step limit in Appendix E, where we observe a slight performance decline and analyze potential causes.
@@ 5.6 Cross-Platform Evaluation
We further evaluated our model on OSWorld to assess cross-platform generalization capabilities. As shown in Table 4, despite being trained exclusively on Windows data, PC Agent-E achieves a 34% relative improvement in Linux systems as well. These results validate the generalizability of our method.
@@
Table 4: Success Rate (%) on OSWorld (30-step).
We also identified an interesting phenomenon in this experiment, which we designate as infeasible hacking in §4: the weaker Qwen2.5-VL-72B model paradoxically achieved markedly better performance on infeasible tasks. This observation suggests that current infeasible task evaluations do not accurately reflect computer use agents’ capabilities. Future research may design better criteria for infeasible tasks, such as checking agents’ rationale when declaring tasks impossible.
6 Conclusion
In this work, we introduce PC Agent-E, an efficient training framework for computer-use agents. Our framework leverages Trajectory Boost, a data synthesis method that augments human demonstrations with diverse action alternatives generated by a frontier model. Trained on as few as 312 augmented trajectories, PC Agent-E achieves a 141% relative improvement over the base model, even outperforms the strong teacher model Claude 3.7 Sonnet. Ablation studies further demonstrate that this approach is not only significantly superior to using human trajectories alone, but also more efficient and effective than direct distillation from the teacher model. These findings suggest that complex computer-use capabilities can be effectively elicited from a remarkably small set of high-quality trajectories.
@@ Appendix A Training Details
Appendix B Evaluation On Original WindowsAgentArena Benchmark
We also evaluate PC Agent-E’s performance on the original WindowsAgentArena benchmark. As shown in Table 5, our model greatly surpasses the previous SOTA method, NAVI, across all task categories. NAVI is a complex agent framework that integrates GPT-4V Achiam et al. (2023) with a specialized tool called Omniparser Lu et al. (2024).
Method Office Web System Coding Media & Video Utils Overall
NAVI Bonatti et al. (2024) 0.0 27.3 33.3 27.3 30.3 8.3 19.5
PC Agent-E (Ours) 2.3 33.1 70.6 37.5 33.3 25.0 27.9
Table 5: Success Rate (%) on the original WindowsAgentArena benchmark.
Appendix C Experiment on Smaller Model
Method Overall
Qwen2.5-VL-7B 5.0
PC Agent-E 7B 6.4
Table 6: Success Rate (%) on WindowsAgentArena-V2.
To test our method on smaller models, we also trained Qwen2.5-VL-7B Bai et al. (2025) on the same 27K dataset for PC Agent-E, resulting in the PC Agent-E 7B. As shown in Table 6, our method yields improvements on the 7B model as well, although the gains are not as significant as those observed with the 72B model. This is because our method mainly enhances the model’s planning ability, as discussed in §5.2, but the small model’s deficiencies in knowledge and grounding limit the overall performance gain.
Appendix D Human Data + Direct Distillation Baseline
[…]
LibreOffice Chrome Edge System VS Code VLC Utils Total
Number of Tasks 42 17 13 24 19 14 12 141
Direct Distillation 2.4 28.8 30.7 50.0 57.9 7.6 25.0 26.2
Human Data + Direct Distillation 2.4 22.9 30.7 54.2 57.9 21.6 16.6 26.9
PC Agent-E (Ours) 4.8 64.1 46.2 50.0 57.9 35.7 33.3 36.0
Table 7: Comparison with Human Data + Direct Distillation baseline on WindowsAgentArena-V2.
Appendix E Extended Test-Time Scaling: 50-Step Evaluation
In addition to the 15-step and 30-step evaluations presented in Figure 8, we conduct an extended evaluation with a 50-step limit to further investigate the test-time scaling behavior of PC Agent-E. As shown in Table 8, both PC Agent-E and the base model Qwen2.5-VL-72B exhibit performance drops when the step limit is extended to 50.
Model 15-step 30-step 50-step
Qwen2.5-VL-72B 11.3 14.9 11.9
PC Agent-E 26.9 36.0 31.4
Table 8: Success Rate (%) on WindowsAgentArena-V2 under varying step limits. Performance improves from 15 to 30 steps but degrades at 50 steps for both models.
Analysis.
We identify that the performance degradation at 50 steps is primarily caused by the models’ inability to correctly recognize task completion. Without a robust stopping mechanism, the additional steps allow the agent to continue acting after a task has already been successfully completed, often leading to actions that undo or disrupt the correct result. We attribute this to the training data distribution (see Table 1 in §3.2), where over 90% of trajectories contain fewer than 15 steps, leaving the model with limited exposure to long-horizon execution and insufficient supervision for task termination.
Scaling within and beyond the training distribution.
Combining the results in Table 8, we observe two distinct regimes. Within a range reasonably close to the training distribution (15 $\rightarrow$ 30 steps), our method demonstrates effective test-time scaling, with performance improving as more steps are allowed. However, when the step limit is pushed significantly beyond the training distribution (50 steps), performance degrades due to the lack of termination awareness. This suggests that the current training paradigm supports near-distribution test-time scaling but does not yet generalize to substantially longer horizons.
Future directions.
We hypothesize that addressing this limitation requires either (1) incorporating longer trajectories into the training data to extend the model’s effective operating range, or (2) introducing targeted supervision for task completion detection, such as an explicit termination prediction objective. We leave the exploration of these directions to future work.
Appendix F Prompts
F.1 Thought Completion